# Invisible Servoing: a Visual Servoing Approach with Return-Conditioned Latent Diffusion

### v2

@@ Abstract
In this paper, we present a novel visual servoing (VS) approach based on latent Denoising Diffusion Probabilistic Models (DDPMs), that explores the application of generative models for vision-based navigation of UAVs (Uncrewed Aerial Vehicles). Opposite to classical VS methods, the proposed approach allows reaching the desired target view, even when the target is initially not visible. This is possible thanks to the learning of a latent representation that the DDPM uses for planning and a dataset of trajectories encompassing target-invisible initial views. A compact representation is learned from raw images using a Cross-Modal Variational Autoencoder. Given the current image, the DDPM generates trajectories in the latent space driving the robotic platform to the desired visual target. The approach has been validated in simulation using two generic multi-rotor UAVs (a quadrotor and a hexarotor). The results show that we can successfully reach the visual target, even if not visible in the initial view.
@@ 1 Introduction
The approach can generalize to different robotic platforms, as the velocity actions planned by the DDPM are tackled by the specific motion controller. We have tested our approach in simulation, applying the planned actions in a receding horizon fashion to allow for online correction of model uncertaintes. Future work includes experiments in real-world and more complex scenarios, and investigation of an end-to-end pipeline.
@@ 4 Methodology
Building upon the planning for control concepts using DDPMs discussed in the previous section, a novel framework for VS is proposed (see Fig. 2). In this approach, the system state is the image frame captured by the robot’s camera. The VS task can be formulated as planning a trajectory of image frames that transitions from the initial frame $\mathbf{x}^{0}$ to the desired target frame $\mathbf{x}^{N}$. This trajectory can be generated using DDPMs by conditioning on the initial and target frames. In this context, the diffusion model fills the gap between the two by generating the intermediate frames. This process is called inpainting [^2] We borrow this terminology from the context of image generation where part of the image pixels is erased and the model is asked to fill in the missing pixels [30].. In this case, the target image is an image of a specific visual feature within the environment that the robot needs to reach. The dataset used to train the diffusion model is a set of different trajectories of image frames captured by the robot’s camera within the environment. The set of images contains frames with and without the visual feature, which is captured from different angles.
@@ 4.1 Dimensionality Reduction using Cross-Modal Variational Autoencoder
The CM-VAE is trained in analogous way to VAE [24] by using a sum of a reconstruction term for each modality and a regularization (KL divergence):
@@ 4.2.1 Return Estimation
In our VS task, the diffusion model is trained on a dataset in which the reward at any step in the trajectory is equal to minus the distance between the visual feature and the robot. Therefore, during planning, if the distance between the visual feature and the robot is known, the return can be estimated by assuming a constant-velocity motion of the robot from the start to the target image. The position vector between the visual feature and the robot at the start and end of the trajectory can be calculated using the feature decoder $p^{\text{f}}_{\boldsymbol{\theta}_{d}}$ of the CM-VAE, which takes the encoded image latent vector z as input and outputs the vector $[r,\theta,\phi,\gamma]$. The spherical coordinates $[r,\theta,\phi]$ can be converted to Cartesian coordinates to get the position vectors $\mathbf{r}_{0}$ and $\mathbf{r}_{N}$ at the start and end of the trajectory respectively, where $N$ is the length of the prediction horizon. The distance vector $\mathbf{d}$ between the start $\mathbf{r}_{0}$ and target position $\mathbf{r}_{N}$ can be calculated as $\mathbf{d}=\mathbf{r}_{0}-\mathbf{r}_{N}$. Assuming that the robot moves with a constant velocity of magnitude $v$ along this vector $\mathbf{d}$ from the start position to the target position, the total travel time can be computedDo we w as $T=||\mathbf{d}||/v$. Therefore, the time between each step $k$ in the trajectory can be calculated as $\Delta t=T/N$. The position vector $\mathbf{r}_{k}$ between the visual feature and the robot at each step $k$ in the trajectory can be computed by sliding along the vector $\mathbf{d}$ with a step size of $v\Delta t$ starting from the start position $\mathbf{r}_{0}$:
@@ 5.1 Return estimation
This analysis clarifies the importance of the return estimation module. To validate it, we verified that the estimated return falls within the interval of admissible values, by doing a sweep analysis in [1.0, 1.0]1.01.0[-1.0,\,1.0][ - 1.0 , 1.0 ] with resolution 0.2 for a given initial and target images. In the challenging scenario where the target is not visible, the estimated return is 0.34, which belongs to the interval [0.2, 0.4]0.20.4[0.2,\,0.4][ 0.2 , 0.4 ], where the two best trajectories were obtained from the sweep analysis.
@@ 6 Conclusion
Future work includes speeding up the sampling process, to allow for a full end-to-end approach. The following strategies will be considered: (i) using more powerful hardware, e.g., Nvidia Jetson, and reducing the diffusion steps using (ii) warm-start strategies, i.e., exploiting previously-generated trajectories as initial guesses for the DDPM to denoise the next trajectory, and (iii) leveraging the research on novel and computationally efficient DDPMs, such as the Denoising Diffusion Implicit Models [32]. Also, the approach will be experimentally validated in real-world and more complex demonstration scenarios.

### v3

# This work was partially funded by the EU: AUTOASSESS project, EU Grant agreement ID: 101120732

@@ Abstract
In this paper, we present a novel visual servoing (VS) approach based on latent Denoising Diffusion Probabilistic Models (DDPMs), that explores the application of generative models for vision-based navigation of UAVs (Uncrewed Aerial Vehicles). Opposite to classical VS methods, the proposed approach allows reaching the desired target view, even when the target is initially not visible. This is possible thanks to the learning of a latent representation that the DDPM uses for planning and a dataset of trajectories encompassing target-invisible initial views. A compact representation is learned from raw images using a Cross-Modal Variational Autoencoder. Given the current image, the DDPM generates trajectories in the latent space driving the robotic platform to the desired visual target. The approach has been validated in simulation using two generic multi-rotor UAVs (a quadrotor and a hexarotor). The results show that we can successfully reach the visual target, even if not visible in the initial view. A video summary with simulations can be found in: https://youtu.be/2Hb3nkkcszE.
@@ 1 Introduction
The approach can generalize to different robotic platforms, as the velocity actions planned by the DDPM are tackled by the specific motion controller. We have tested our approach in simulation, applying the planned actions in a receding horizon fashion to allow for online correction of model uncertainties. Future work includes experiments in real-world and more complex scenarios, and investigation of an end-to-end pipeline.
@@ 4 Methodology
Building upon the planning for control concepts using DDPMs discussed in the previous section, a novel framework for VS is proposed (see Fig. 2). In this approach, the system state is the image frame captured by the robot’s camera. The VS task can be formulated as planning a trajectory of image frames that transitions from the initial frame $\mathbf{x}^{0}$ to the desired target frame $\mathbf{x}^{N}$. This trajectory can be generated using DDPMs by conditioning on the initial and target frames. In this context, the diffusion model fills the gap between the two by generating the intermediate frames. This process is called inpainting [^3] We borrow this terminology from the context of image generation where part of the image pixels is erased and the model is asked to fill in the missing pixels [30].. In this case, the target image is an image of a specific visual feature within the environment that the robot needs to reach. The dataset used to train the diffusion model is a set of different trajectories of image frames captured by the robot’s camera within the environment. Each trajectory is generated by randomly spawning the robot at a random position and orientation in the environment and then tasking it to reach and observe a specified target. The set of images contains frames with and without the visual feature, which is captured from different angles.
@@ 4.1 Dimensionality Reduction using Cross-Modal Variational Autoencoder
The CM-VAE is trained in an analogous way to a VAE [24] by maximizing the log likelihood (reconstruction terms) of image $\mathbf{x}_{\text{rgb}}$ and coordinate frame $\mathbf{x}_{\text{f}}$ given the latent vector $\mathbf{z}$ and $\tilde{\mathbf{z}}$, respectively, while minimizing the KL divergence (regularization term) between the encoder’s output distribution $q_{\text{rgb}}(\textbf{z}|\textbf{x}_{\text{rgb}})$ and a standard normal distribution $\mathcal{N}(\mathbf{0},I)$, representing the prior distribution of the latent vectors $p(\mathbf{z})$:
@@ 4.2.1 Return Estimation
In our VS task, the diffusion model is trained on a dataset in which the reward at any step in the trajectory is equal to minus the distance between the visual feature and the robot. Therefore, during planning, if the distance between the visual feature and the robot is known, the return can be estimated by assuming a constant-velocity motion of the robot from the start to the target image. The position vector between the visual feature and the robot at the start and end of the trajectory can be calculated using the feature decoder $p^{\text{f}}_{\boldsymbol{\theta}_{d}}$ of the CM-VAE, which takes the encoded image latent vector z as input and outputs the vector $[r,\theta,\phi,\gamma]$. The spherical coordinates $[r,\theta,\phi]$ can be converted to Cartesian coordinates to get the position vectors $\mathbf{r}_{0}$ and $\mathbf{r}_{N}$ at the start and end of the trajectory respectively, where $N$ is the length of the prediction horizon. The distance vector $\mathbf{d}$ between the start $\mathbf{r}_{0}$ and target position $\mathbf{r}_{N}$ can be calculated as $\mathbf{d}=\mathbf{r}_{0}-\mathbf{r}_{N}$. Assuming that the robot moves with a constant velocity of magnitude $v$ along this vector $\mathbf{d}$ from the start position to the target position, the total travel time can be computed as $T=||\mathbf{d}||/v$. Therefore, the time between each step $k$ in the trajectory can be calculated as $\Delta t=T/N$. The position vector $\mathbf{r}_{k}$ between the visual feature and the robot at each step $k$ in the trajectory can be computed by sliding along the vector $\mathbf{d}$ with a step size of $v\Delta t$ starting from the start position $\mathbf{r}_{0}$:
@@ 5.1 Return estimation
This analysis clarifies the importance of the return estimation module. To validate it, we verified that the estimated return falls within the interval of admissible values, by doing a sweep analysis in [1.0, 1.0]1.01.0[-1.0,\,1.0][ - 1.0 , 1.0 ] with resolution 0.2 for a given initial and target images. In a challenging scenario where the target is not visible, a sweep analysis has shown that the interval of admissible values resulting in a planned trajectory that smoothly converges to the target image without abrupt jumps is [0.2, 0.4]0.20.4[0.2,\,0.4][ 0.2 , 0.4 ]. The return estimated by the algorithm described in Sec. 4.2.1 is 0.34, which is consistent with the result of the sweep analysis.
@@ 6 Conclusion
In the current implementation, the trajectory generated by the DDPM is followed by relying on a Motion Capture system for localization. As future work, we aim to remove the dependency on a localization pipeline by developing a full end-to-end approach. This way, the invisible servoing algorithm will overcome the limitations of traditional localization pipelines, which struggle in case of lack of features or bad lighting. Indeed, such challenging conditions can be reproduced in the dataset and the models trained to guarantee successful servoing. To achieve the objective of an end-to-end pipeline, the sampling process needs to be sped up. The following strategies will be considered: (i) using more powerful hardware, e.g., Nvidia Jetson, and reducing the diffusion steps using (ii) warm-start strategies, i.e., exploiting previously-generated trajectories as initial guesses for the DDPM to denoise the next trajectory, and (iii) leveraging the research on novel and computationally efficient DDPMs, such as the Denoising Diffusion Implicit Models [32]. Future work also includes combining the DDMP approach with traditional MPC. This would allow including additional objectives and dynamic feasibility of the solution.
In addition, the approach will be experimentally validated in real-world and more complex demonstration scenarios.